# Quantum-Assisted Vehicle Routing: Realizing QAOA-based Approach on Gate-Based Quantum Computer

### v1

@@
Abstract
The Vehicle Routing Problem (VRP) is a crucial optimization challenge with significant economic and environmental implications, particularly in logistics and transportation planning. While classical algorithms struggle to efficiently solve large-scale instances of VRP due to its combinatorial complexity, quantum computing presents a promising alternative for tackling such problems. In this work, we explore the application of the Quantum Approximate Optimization Algorithm (QAOA) to solve instances of VRP, analyzing its effectiveness and scalability. We formulate VRP as a Quadratic Unconstrained Binary Optimization (QUBO) problem by encoding the constraints into a single cost function suitable for QAOA. Our study investigates the impact of problem size on quantum circuit complexity and evaluate the feasibility of executing QAOA-based VRP solutions on near-term quantum hardware. The results indicate that while QAOA demonstrates potential for solving VRP, the primary limitation lies in circuit depth and noise-induced errors, which critically affect performance on current quantum processors. Overcoming these challenges will require advancements in error mitigation techniques and more efficient quantum circuit designs to realize the full potential of quantum computing for combinatorial optimization.
Index Terms:
@@ Abstract
Quantum computing, Vehicle routing problem, Quantum Approximate Optimization Algorithm, Gate-based quantum computer, Quadratic unconstrained binary optimization
I Introduction
Metropolitan areas serve as dense hubs of economic activity, contributing over 60% of global GDP while also being responsible for approximately 70% of fossil fuel-related CO2 emissions [1]. These contributions highlight a critical challenge, that while cities drive economic growth, they also create substantial traffic congestion and environmental externalities. The role of urban logistics management and optimization can play a crucial role in addressing the inefficiencies while ensuring economic and environmental sustainability [2]. However, this optimization is complex due to the vast number of variables, constraints and interactions. At the heart of it lies the Vehicle Routing Problem (VRP), a fundamental combinatorial optimization problem that seeks to determine the most efficient routes for a fleet of vehicles to serve a given set of customers while minimizing operational cost such as the distance traveled, fuel consumption, or delivery times [3]. It is an extension of the well known NP-hard Traveling Salesman Problem (TSP) [4], and is widely used in logistics, supply chain management, and transportation planning, with applications ranging from last-mile delivery to large-scale freight distribution.
Since classical computers struggle to efficiently solve large-scale NP-hard problems due to their exponential computational complexity, alternative computing paradigms may offer new pathways to tackle these challenges. Quantum computing is a rapidly advancing field that harnesses the principles of quantum mechanics to solve problems that are intractable for classical computers. Unlike traditional digital computing, which relies on bits that assume values of either 0 or 1, quantum computing uses quantum bits (qubits) that can exist in a superposition of both 0 and 1. Additionally, quantum computers utilize entanglement, a phenomenon where the state of one qubit is intrinsically linked to another, which enables highly correlated operations that surpass classical information processing capabilities. The field is progressing swiftly in this noisy intermediate-scale quantum (NISQ) era, characterized by sub-1000 qubit quantum processors, where noise and decoherence present significant challenges [5]. Despite these limitations, ongoing advancements in hardware, error mitigation techniques, and quantum algorithms continue to push the boundaries of what is computationally feasible, paving the way for practical quantum advantage.
@@ I Introduction
Figure 1: Solving the VRP using QAOA involves a formulation of the link based constrained minimization form to the unconstrained quadratic binary representation, which is then cast as an Ising Hamiltonian required for QAOA. The high level circuit is transpiled into one and two qubit basic gates for implementation on the IBM-Rensselaer quantum computer. The circuit is run repeatedly to optimize the variational parameters, and a final sampling provides the solution. The problem graph shown here is for illustrative purposes only. Section IV describes the specific problem studied.
Quantum computational models can broadly be categorized into gate-based quantum computing, quantum annealing, and other specialized paradigms such as continuous-variable quantum computing [6] and topological quantum computing, each optimized for distinct computational tasks and problem domains. Special purpose quantum devices such as quantum annealers are able to solve combinatorial optimization problems by using a continuous quantum adiabatic evolution process to find low energy states of a given problem [7]. They rely on quantum tunneling through energy barriers, which allows the state to escape local minima. The fundamental concept is that the system starts in a known ground state and evolves slowly to an unknown ground state of the problem Hamiltonian, which represents the optimization objective.
Gate-based quantum computers operate on the quantum circuit model, where computations are performed using quantum gates, i.e., unitary operations that manipulate qubits similarly to how logic gates operate on classical bits. Algorithms such as Shor’s factoring algorithm and Grover’s search algorithm rely on a series of controlled quantum gate operations to achieve computational speedups over their classical counterparts. Gate-based quantum computers are also called universal in the sense that they can implement any quantum algorithm up to a reasonable approximation [8]. This gives them much more flexibility, but requires significant error correction to scale up.
One promising method of solving optimization problems on gate based quantum computers is the Quantum Approximate Optimization Algorithm (QAOA) [9], inspired by the quantum adiabatic theorem that is fundamental to quantum annealing [10]. QAOA is a hybrid algorithm using both quantum and classical hardware to approximate solutions to combinatorial optimization problems. It encodes the problem into a quantum circuit, applying a sequence of alternating unitary operators, and optimizing parameters to maximize the probability of measuring a good solution. An $n$-qubit QAOA circuit can encode and process a solution space of size $2^{n}$, providing near optimal solutions in polynomial time for the maximum cut problem in graph theory [11]. Due to its suitability for near-term quantum devices, QAOA has gained attention for tackling problems in logistics, finance, and network optimization.
In this paper we explore the application of QAOA to the VRP, using quantum computing to tackle its inherent complexity. While the problem is NP-hard, the inherent graph structure of VRP makes it a compelling candidate for quantum approaches. We formulate VRP as a Quadratic Unconstrained Binary Optimization (QUBO) problem to implement QAOA and analyze its performance while highlighting the challenges associated with the approach. To the best of our knowledge, while there have been studies on using quantum simulation for VRP, this paper is among the first efforts to solve the VRP using QAOA implemented on a gate-based quantum computer; we also considered subtour constraints in the modeling and quantum implementation process. In addition, we discuss key methodological considerations, including problem encoding, circuit depth, parameter optimization, and hardware-induced limitations such as noise and decoherence. These findings provide insights into both the current limitations and potential avenues of further investigation in quantum optimization techniques for solving real-world routing problems.
II Literature Review
@@
A universal quantum computer refers to a system that can implement any quantum algorithm and perform arbitrary quantum computations similar to how a Turing machine can execute any classical algorithm, an idea put forth by Richard Feynman [12]. A gate based quantum computer can simulate any unitary evolution efficiently using a set of universal quantum gates capable of approximating the unitary operation to arbitrary precision. The theoretical foundations for quantum computation were laid down in [8] showing that certain tasks involving randomness can be performed by a quantum computer faster than classical probabilistic algorithms on a classical computer. Some examples of known quantum advantage are factorization using Shor’s algorithm [13], unstructured search using Grover’s algorithm [14], and simulation of the electronic structures of molecules or other quantum systems [15].
@@
Special-purpose quantum computers such as quantum annealers are built to exploit quantum mechanics for specific problem domains. Similar to thermal simulated annealing [16], quantum annealing is a process that finds the lowest-energy configuration of a system by gradually transforming from a simple quantum state to one representing the optimal solution [17]. It was shown to outperform classical methods for the traveling salesperson problem [18], and has since found commercial use through companies such as D-Wave that are now offering devices with over 5000 qubits which see use in logistics, finance, and machine learning [19]. However, special-purpose computers exhibit a speedup that is problem-dependent, and can only perform a narrow set of tasks and lack the flexibility to run any general quantum algorithm [20]. The capacitated vehicle routing problem has been approached by a hybrid algorithm utilizing quantum annealing, with the results showing no clear benefit over classical solvers [21].
@@
A Variational Quantum Algorithm (VQA) is a hybrid quantum-classical approach for universal quantum computing that leverages quantum circuits parameterized by classical variables to solve optimization problems. For example, the Variational Quantum Eigensolver (VQE) seeks to find the minimum eigenvalue or the ground state energy of a Hamiltonian, and has seen application in materials science and quantum chemistry [22]. Another VQA, the Quantum Approximate Optimization Algorithm (QAOA) was introduced for approximate combinatorial optimization that aims to find the bit string which maximizes a cost function [9]. It guarantees the optimization in the limit of infinite layers in a manner similar to the adiabatic evolution used in quantum annealing [10].
@@
Beyond its use for solving optimization problems, QAOA may be a route to establishing quantum supremacy since classically emulating QAOA in an efficient way would collapse the polynomial hierarchy of computational complexity [11]. At low depth proportional to $\log n$, the performance of QAOA is known to be low as the algorithm does not see the whole graph in Max-Cut type problems [23]. There are diminishing returns towards accuracy versus QAOA circuit depth [24]. In fact, a study into the classical optimizers and circuit depth on noisy quantum devices concluded that beyond 6 to 8 layers the probability of success decreases due to noise accumulation and decoherence [25]. The upper bound on QAOA is yet to be determined in practical scenarios but it has been observed that it outperformed the best classical solvers for large boolean satisfiability problems [26]. Importantly, using the QAOA outputs as guesses to the classical optimizer did not yield a significantly better overall performance.
@@
While interest in quantum computing for intelligent transportation systems has been expressed only recently [27][28], the combinatorial optimization problems of TSP and VRP, have received significant attention from a quantum computing perspective [18] [29]. Classical methods exist to solve the problem exactly, for example branch and bound [30], which scale exponentially with problem size. On the other hand metaheuristic algorithms can provide approximate solutions, like Tabu search [31] and genetic algorithms [32], but they are not guaranteed to find the exact optimal solution in large problems. QAOA was applied to small scale TSP instances, giving high quality solutions at limited depth with a mixer Hamiltonian corresponding to the problem constraints [33]. The performance of QAOA simulation on a VRP formulation without subtour elimination constraints was found to depend strongly on the classical optimization method and the depth of the ansatz since training time can increase exponentially, but this may be prevented by a good initialization [34]. Another study discovered that QAOA biases solutions toward lower energies but fails to distinguish between feasible solutions of different costs, limiting its effectiveness for VRP [35]. The authors conjecture that this may be due to a small energy gap in the Hamiltonian representation, which could be mitigated by reformulating the problem to increase energy separation. A detailed comparison of different VRP formulations and quantum algorithms demonstrated that QAOA with constrained optimization by linear approximation (COBYLA) delivered the highest probability of sampling the optimal configuration given enough post optimization sampling shots [36].
III Conceptual Overview
III-A Quantum Computing
@@
$|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$ (1)
[…]
The state of multiple qubits lies in an exponential state space. If there are $n$ qubits, each of the $2^{n}$ basis states has a complex coefficient, $\alpha_{i}$, that dictates the probability, $|\alpha_{i}|^{2}$, of measuring that basis state. It is for this reason that quantum computing cannot be simulated on classical computers [37]. For example, 50 qubits will require $2^{50}$ complex numbers to describe the state, and any circuit operations will require multiplication with a matrix of size $2^{50}\times 2^{50}$. If each coefficient was represented with just one byte, the state vector would take over 1000 Terabytes to store in a classical computer.
The IBM Eagle type quantum computer uses transmon qubits which are a type of superconducting charge qubit, similar to the Cooper-pair box. At a fundamental level the transmon qubit works on the principle of the Josephson effect [38], which is the appearance of a constant current through an insulating barrier between two superconductors in the absence of an external electromagnetic field due to quantum tunneling. This effect enables quantum coherence, and combined with precise microwave control, makes the transmon a stable and controllable qubit for quantum computing [39]. Microwave radiation can induce quantized voltages across the junction, allowing control, coupling, and measurement using microwave resonators. The qubit states $|0\rangle$ and $|1\rangle$ correspond to the two lowest energy levels of the transmon. There also exist physical coupling channels on the quantum device that allow direct capacitive coupling between qubits to implement two-qubit entangling gates.
@@
where $i=\sqrt{-1}$, $\hslash$ is the reduced Planck’s constant, and $\mathbf{H}^{2^{n}\times 2^{n}}$ is a Hermitian matrix known as the Hamiltonian. The solution of the equation yields an exponential operator $U(t)=e^{i\mathbf{H}t/\hslash}$, for any $t>0$,
@@
These operations can transform a known quantum state to any desired state in a quantum computer. One basic gate is the Hadamard gate, which, when applied to a qubit in the zero state, converts it into an even superposition of basis states $|0\rangle$ and $|1\rangle$.
@@
Applying it to $|1\rangle$ creates an even superposition as well, but with a phase difference. This phase difference is also a feature of quantum states not found in classical stochastic systems.
@@
Figure 2: Two states showing even superposition on the Bloch sphere. The Z direction is upwards.
If measured on the computational basis $|0\rangle,|1\rangle$ on a quantum computer, the samples of 0s and 1s from these states would be equal, subject to noise. The state of a qubit can also be depicted on the Bloch spheres in Figure 2, which show the two Hadamard states pointing to opposite points on the equator, each representing an even measurement probability between the two poles of the computational basis on the Z axis. The Bloch sphere representation can be derived from Equation 1,
@@
Figure 3: Quantum circuit for creating the Bell state. Qubits start in the $|00\rangle$ state. The element H is a Hadamard gate applied to qubit-0, while the blue symbol with $+$ sign indicates a CNOT gate, with $q_{0}$ controlling the NOT operation on $q_{1}$. The two gray elements on the right depict measurement.
@@
$\begin{split}|\phi\rangle=H\otimes I|00\rangle=\frac{1}{\sqrt{2}}\begin{bmatrix}1&0&1&0\\ 0&1&0&1\\ 1&0&-1&0\\ 0&1&0&-1\end{bmatrix}\begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix}\\ |\phi\rangle=\frac{1}{\sqrt{2}}\begin{bmatrix}1\\ 0\\ 1\\ 0\end{bmatrix}=\frac{1}{\sqrt{2}}\begin{bmatrix}1\\ 0\\ 0\\ 0\end{bmatrix}+0\begin{bmatrix}0\\ 1\\ 0\\ 0\end{bmatrix}+\frac{1}{\sqrt{2}}\begin{bmatrix}0\\ 0\\ 1\\ 0\end{bmatrix}+0\begin{bmatrix}0\\ 0\\ 0\\ 1\end{bmatrix}\\ \Rightarrow\ |\phi\rangle=\frac{1}{\sqrt{2}}\left(|00\rangle+|10\rangle\right)\\ \end{split}$

### v2

@@
Abstract.
The Vehicle Routing Problem (VRP) is a fundamental combinatorial optimization challenge with broad applications in logistics and transportation. In this work, we present a quantum-assisted framework that integrates the Quantum Approximate Optimization Algorithm (QAOA) with a link-based formulation of VRP. Our approach encodes flow conservation and subtour elimination directly into the cost Hamiltonian, preserving graph structure while avoiding ancillary qubits. We design and implement the full pipeline on a gate-based quantum computer, including problem encoding, circuit synthesis, and execution on IBM Quantum System One. Experimental results on small VRP instances highlight the effects of penalty scaling, coefficient normalization, and circuit depth on solution feasibility under hardware noise. While scalability remains constrained by circuit complexity and decoherence, the study demonstrates a practical pathway for implementing VRP on quantum hardware and identifies methodological directions for advancing near-term quantum optimization.
@@ Abstract.
Quantum computing, Vehicle routing problem, Quantum Approximate Optimization Algorithm, Quadratic unconstrained binary optimization
journal: JATSjournalvolume: 37journalnumber: 4article: 111publicationmonth: 11ccs: Computer systems organization Quantum computingccs: General and reference Experimentationccs: Mathematics of computing Combinatorial optimizationcopyright: acmlicensedjournal: TQCjournalyear: 2025journalvolume: 1journalnumber: 1article: 1publicationmonth: 1doi: 10.1145/0000003
1. Introduction
Metropolitan areas generate over 60% of global GDP but account for nearly 70% of fossil fuel-related CO2 emissions (NASA, 2024), underscoring the dual challenge of economic growth and environmental externalities. Urban logistics optimization is central to mitigating congestion and inefficiencies while promoting sustainability (He et al., 2023), yet remains complex due to numerous variables and constraints. At its core lies the Vehicle Routing Problem (VRP, a fundamental NP-hard extension of the Traveling Salesman Problem (TSP) (Clarke and Wright, 1964; Jünger et al., 1995), which seeks efficient vehicle routes to minimize costs such as distance, fuel, or time, and has broad applications from last mile delivery to large scale freight distribution.
VRP instances arise in logistics and transportation planning, from last-mile delivery to large-scale distribution, and even modest-sized instances are NP-hard and computationally intractable to solve optimally with classical algorithms. This motivates exploring alternative computing paradigms for VRP and similar combinatorial problems. In particular, quantum computing has emerged as a promising paradigm for tackling NP-hard problems by exploiting quantum superposition and entanglement to potentially search huge solution spaces more efficiently. Although current quantum devices operate in the noisy intermediate-scale quantum (NISQ) regime with only tens to hundreds of qubits and significant noise, the field is advancing rapidly (Preskill, 2018). Quantum algorithms for optimization, if carefully designed to work within these hardware limits, could eventually augment classical high-performance computing (HPC) in solving real-world optimization tasks.
@@ 1. Introduction
Figure 1. Solving the VRP using QAOA involves a formulation of the link based constrained minimization form to the unconstrained quadratic binary representation, which is then cast as an Ising Hamiltonian required for QAOA. The high level circuit is transpiled into one and two qubit basic gates for implementation on the IBM-Rensselaer quantum computer. The circuit is run repeatedly to optimize the variational parameters, and a final sampling provides the solution. The problem graph shown here is for illustrative purposes only. Section IV describes the specific problem studied.
One leading approach for combinatorial optimization on gate-based quantum computers is the Quantum Approximate Optimization Algorithm (QAOA) (Farhi et al., 2014), inspired by the quantum adiabatic theorem that is fundamental to quantum annealing (Farhi et al., 2000). QAOA is a hybrid quantum-classical algorithm that encodes an optimization problem into a parameterized quantum circuit and uses a classical optimizer to tune those parameters. By alternating problem specific cost operators and mixing operators on $n$ qubits, QAOA can in principle concentrate probability mass on high quality solutions among the $2^{n}$ possible states. Its suitability for near-term devices has made QAOA a focal point for studying problems in logistics, finance, and network optimization. Applying QAOA to the VRP is especially enticing because VRP’s graph-structured decision space might be naturally encoded in a quantum Hamiltonian. A quantum enhanced solution to VRP could not only demonstrate the capabilities of NISQ algorithms on a problem of practical relevance, but also pave the way for quantum assisted HPC workflows where classical and quantum resources are integrated to tackle industry-scale optimization (Elsharkawy et al., 2025).
However, significant methodological challenges arise in bringing QAOA to VRP. First, one must devise an encoding of VRP into a quantum-friendly cost function (e.g. a Quadratic Unconstrained Binary Optimization, QUBO) that captures all the problem constraints without exploding the resource requirements. The VRP has complex feasibility constraints: each customer must be visited exactly once, vehicles have route length or capacity limits, and routes must form contiguous tours, which are nontrivial to represent in a qubit Hamiltonian. Naively encoding all constraints can require a large number of qubits or high order penalty terms. Prior attempts have taken different formulation strategies, for example, some works introduced time step indexed decision variables to linearize the route order, thereby restricting the search to classically feasible sequences (Palackal et al., 2023). This approach can asymptotically reduce the quantum resources needed by eliminating invalid states at the cost of additional ancilla qubits and complex multi-controlled operations. Another approach is to use unbalanced penalty weights for soft constraints, effectively biasing the cost function to discourage violations without strictly forbidding them; but this requires careful tuning of penalty coefficients via classical optimization and can introduce higher order interactions (Montañez-Barrera et al., 2024). Designing an encoding that is compact yet enforces VRP constraints remains a core challenge.
Ensuring solution feasibility is a key challenge when applying QAOA to constrained problems. Since the quantum state explores superpositions of many bitstrings, an unsuitable Hamiltonian or mixer can increase the likelihood of measuring infeasible routes (e.g., routes omitting customers or forming subtours). To address this, researchers have explored customized mixers that restrict the evolution to feasible subspaces, as well as warm-start strategies that initialize QAOA in valid solutions. However, empirical studies on VRP have shown only limited improvements from these techniques (Fitzek et al., 2024). For example, even in hybrid approaches where classical clustering was used to reduce the problem before applying QAOA, performance remained behind comparable VQE methods, with QAOA struggling to consistently recover feasible solutions in small instances. These findings highlight the ongoing difficulty of steering QAOA effectively toward the feasible region of combinatorial problems.
Furthermore, quantum circuit depth and hardware limitations impose severe restrictions on solving VRP with QAOA. Each layer of QAOA applies numerous two qubit gates to implement problem interactions. As the problem size or number of QAOA layers increases, the circuit quickly becomes deeper and more error-prone. Current superconducting quantum processors have limited qubit connectivity and coherence times, meaning complex QAOA circuits suffer from accumulated gate errors and decoherence before completion (Blekos et al., 2024). For a VRP instance of even modest size, achieving good solution quality might require many layers of QAOA, a regime simply not attainable on today’s hardware. Moreover, qubit count and topology are limiting factors: encoding a VRP with $N$ locations typically involves $O(N^{2})$ binary decision variables (for example, representing each possible directed edge or each assignment of a location to a route position). This can easily exceed the number of physical qubits available, and even when it does not, mapping the logical interactions onto a real device (such as IBM’s heavy-hex lattice) can require additional swap gates that further increase the effective circuit depth.
In this work, we go beyond a direct application of QAOA to VRP by co-designing quantum circuits with the problem’s graph structure. We develop both the algorithmic formulation and its implementation framework on gate-based quantum hardware, i.e., IBM Quantum System One. While the problem is NP-hard, the inherent graph structure of VRP makes it a compelling candidate for quantum approaches. Unlike approaches that precalculate the clusters on a classical computer (Feld et al., 2019), or use time expanded versions of the VRP to trade off qubits for constraint simplification (Palackal et al., 2023), we design the quantum solution based on a link based optimization formulation. Using directional links as decision variables we naturally capture the graph structure of VRP and this allows us to incorporate subtour elimination constraints and flow conservation directly into the cost Hamiltonian while avoiding ancillary qubits.
To the best of our knowledge, while there have been studies on using quantum simulation for VRP, this paper is among the first efforts to design a comprehensive pipeline from modeling to implementing VRP on a gate-based quantum computer. Moreover, the study examines key methodological aspects, including problem encoding, circuit depth, parameter optimization, and hardware related limitations such as noise and decoherence. We found that the choice of penalty scaling and coefficient normalization significantly affects solution feasibility in QAOA. In particular, normalization of Hamiltonian terms improved performance on larger VRP instances, while penalty values around twice the sum of edge weights provided consistently better feasibility. Together, these analyses shed light on the present challenges and point to promising directions for advancing quantum optimization techniques in real-world routing applications.
2. Literature Review
@@
A universal quantum computer refers to a system that can implement any quantum algorithm and perform arbitrary quantum computations similar to how a Turing machine can execute any classical algorithm, an idea put forth by Richard Feynman (Feynman, 1982). A gate based quantum computer can simulate any unitary evolution efficiently using a set of universal quantum gates capable of approximating the unitary operation to arbitrary precision. The theoretical foundations for quantum computation were laid down in (Deutsch, 1985) showing that certain tasks involving randomness can be performed by a quantum computer faster than classical probabilistic algorithms on a classical computer. Some examples of known quantum advantage are factorization using Shor’s algorithm (Shor, 1999), unstructured search using Grover’s algorithm (Grover, 1996), and simulation of the electronic structures of molecules or other quantum systems (Tacchino et al., 2020).
@@
Special-purpose quantum computers such as quantum annealers are built to exploit quantum mechanics for specific problem domains. Similar to thermal simulated annealing (Kirkpatrick et al., 1983), quantum annealing is a process that finds the lowest-energy configuration of a system by gradually transforming from a simple quantum state to one representing the optimal solution (Finnila et al., 1994). It was shown to outperform classical methods for the traveling salesperson problem (Martoňák et al., 2004), and has since found commercial use through companies such as D-Wave that are now offering devices with over 5000 qubits which see use in logistics, finance, and machine learning (Boothby et al., 2020). However, special-purpose computers exhibit a speedup that is problem-dependent, and can only perform a narrow set of tasks and lack the flexibility to run any general quantum algorithm (Yarkoni et al., 2022). The capacitated vehicle routing problem has been approached by a hybrid algorithm utilizing quantum annealing, with the results showing no clear benefit over classical solvers (Feld et al., 2019).
@@
A Variational Quantum Algorithm (VQA) is a hybrid quantum-classical approach for universal quantum computing that leverages quantum circuits parameterized by classical variables to solve optimization problems. For example, the Variational Quantum Eigensolver (VQE) seeks to find the minimum eigenvalue or the ground state energy of a Hamiltonian, and has seen application in materials science and quantum chemistry (McClean et al., 2016). Another VQA, the Quantum Approximate Optimization Algorithm (QAOA) was introduced for approximate combinatorial optimization that aims to find the bit string which maximizes a cost function (Farhi et al., 2014). It guarantees the optimization in the limit of infinite layers in a manner similar to the adiabatic evolution used in quantum annealing (Farhi et al., 2000).
@@
Beyond its use for solving optimization problems, QAOA may be a route to study establishing quantum supremacy since classically emulating QAOA in an efficient way would collapse the polynomial hierarchy of computational complexity. However this is believed to be unlikely (Farhi and Harrow, 2016). At low depth proportional to $\log n$, the performance of QAOA is known to be low as the algorithm does not see the whole graph in Max-Cut type problems (Farhi et al., 2020). There are diminishing returns towards accuracy versus QAOA circuit depth (Blekos et al., 2024). In fact, a study into the classical optimizers and circuit depth on noisy quantum devices concluded that beyond 6 to 8 layers the probability of success decreases due to noise accumulation and decoherence (Pellow-Jarman et al., 2024). The upper bound on QAOA is yet to be determined in practical scenarios but it has been observed that it outperformed the best classical solvers for large boolean satisfiability problems (Boulebnane and Montanaro, 2024). Importantly, using the QAOA outputs as guesses to the classical optimizer did not yield a significantly better overall performance.
@@
While interest in quantum computing for intelligent transportation systems has been expressed only recently (Burkacky et al., 2020)(Zhuang et al., 2024), the combinatorial optimization problems of TSP and VRP, have received significant attention from a quantum computing perspective (Martoňák et al., 2004)(Leonidas et al., 2023). Classical methods exist to solve the problem exactly, for example branch and bound (Theurich et al., 2021), which scale exponentially with problem size. On the other hand metaheuristic algorithms can provide approximate solutions, like Tabu search (Sicilia et al., 2016) and genetic algorithms (Baker and Ayechew, 2003), but they are not guaranteed to find the exact optimal solution in large problems. QAOA was applied to small scale TSP instances, giving high quality solutions at limited depth with a mixer Hamiltonian corresponding to the problem constraints (Ruan et al., 2020). The performance of QAOA simulation on a VRP formulation without subtour elimination constraints was found to depend strongly on the classical optimization method and the depth of the ansatz since training time can increase exponentially, but this may be prevented by a good initialization (Azad et al., 2022). Another study discovered that QAOA biases solutions toward lower energies but fails to distinguish between feasible solutions of different costs, limiting its effectiveness for VRP (Fitzek et al., 2024). The authors conjecture that this may be due to a small energy gap in the Hamiltonian representation, which could be mitigated by reformulating the problem to increase energy separation. A detailed comparison of different VRP formulations and quantum algorithms demonstrated that QAOA with constrained optimization by linear approximation (COBYLA) delivered the highest probability of sampling the optimal configuration given enough post optimization sampling shots (Harwood et al., 2021).
Using a hybrid approach of clustering on a classical computer and then solving the subproblem of capacitated VRP using QAOA and VQE was tested in (Palackal et al., 2023). Despite the time expanded formulation, VQE achieved feasible solutions on small instances, QAOA consistently underperformed, and extensions such as constraint-preserving mixers and warm-start QAOA did not improve results. Similar approaches that incorporate time steps into the decision variables and restrict the state space to feasible bitstrings achieve asymptotically lower resource requirements, but in practice incur substantial ancilla qubit overhead for small instances (Xie et al., 2024). Other approaches to simplify the quantum circuit involve the use of unbalanced penalization for inequality constraints (Montañez-Barrera et al., 2024), however, these require classical optimization to determine the penalty coefficients.
In this work, we formulate the VRP using directional links as decision variables with subtour elimination constraints, and investigate the implementation of QAOA, focusing on the encoding strategy, circuit depth, and scalability on a real quantum computer.
3. Conceptual Overview
3.1. Quantum Computing
@@
(1) $|\psi\rangle=\alpha|0\rangle+\beta|1\rangle$
where $\alpha$ and $\beta$ are probability amplitudes, complex numbers satisfying the completeness relation $|\alpha|^{2}+|\beta|^{2}=1$. The notation $|\cdot\rangle$ is called a ‘ket’ which can be represented as a column vector, so $|0\rangle=\begin{bmatrix}1\\ 0\end{bmatrix}$, and $|1\rangle=\begin{bmatrix}0\\ 1\end{bmatrix}$. The ‘bra’ notation, $\langle\cdot|$, is the adjoint of ket and can be represented as a row vector.
The state of multiple qubits lies in an exponential state space. If there are $n$ qubits, each of the $2^{n}$ basis states has a complex coefficient, $\alpha_{i}$, that dictates the probability, $|\alpha_{i}|^{2}$, of measuring that basis state. It is for this reason that quantum computing cannot be efficiently simulated on classical computers (Bremner et al., 2011). For example, 50 qubits will require $2^{50}$ complex numbers to describe the state, and any circuit operations will require multiplication with a matrix of size $2^{50}\times 2^{50}$. If each coefficient was represented with just one byte, the state vector would take over 1000 Terabytes to store in a classical computer.
The IBM Eagle type quantum computer uses transmon qubits which are a type of superconducting charge qubit, similar to the Cooper-pair box. At a fundamental level the transmon qubit works on the principle of the Josephson effect (Josephson, 1964), which is the appearance of a constant current through an insulating barrier between two superconductors in the absence of an external electromagnetic field due to quantum tunneling. This effect enables quantum coherence, and combined with precise microwave control, makes the transmon a stable and controllable qubit for quantum computing (Koch et al., 2007). Microwave radiation can induce quantized voltages across the junction, allowing control, coupling, and measurement using microwave resonators. The qubit states $|0\rangle$ and $|1\rangle$ are typically assigned to the two lowest energy levels of the transmon. There also exist physical coupling channels on the quantum device that allow direct capacitive coupling between qubits to implement two-qubit entangling gates.
@@
[…]
@@
These operations can transform a known quantum state to any desired state in a quantum computer. One basic gate is the Hadamard gate, which, when applied to a qubit in the zero state, converts it into an equal superposition of basis states $|0\rangle$ and $|1\rangle$.
@@
Applying it to $|1\rangle$ creates an equal superposition as well, but the signs of the probability amplitudes differ. This characteristic of probability amplitudes allows quantum states to interfere, which is not possible in classical stochastic systems which are described using probabilities.
@@
Figure 2. Two states showing even superposition on the Bloch sphere. The Z direction is upwards.
If measured on the computational basis $|0\rangle,|1\rangle$ on a quantum computer, the samples of 0s and 1s from these states would be equal, subject to noise. The state of a qubit can also be depicted on the Bloch spheres in Figure 2, which show the two Hadamard states pointing to opposite points on the equator, each representing an equal measurement probability between the two poles of the computational basis on the Z axis. The Bloch sphere representation can be derived from Equation 1,
@@
Figure 3. Quantum circuit for creating the Bell state. Qubits start in the $|00\rangle$ state. The element H is a Hadamard gate applied to qubit-0, while the blue symbol with ++ sign indicates a CNOT gate, with $q_{0}$ controlling the NOT operation on $q_{1}$. The two gray elements on the right depict measurement.
@@
[…]